OpenStudy (anonymous):

8th grade math help please? I will medal :D I just have 5 questions...

OpenStudy (anonymous):

lets see them

OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? x+y=-1 x-3y=11 A. (2, −3) B. (−2, 1) C. (5, −2) D. (−6, 5)

OpenStudy (anonymous):

I will post them one at a time, posting them all at once makes it hard to understand

OpenStudy (anonymous):

@adelyn_helps XD I tagged you

OpenStudy (adelyn_helps):

I am here :D

OpenStudy (anonymous):

do you know this...? this is hard for me :(

OpenStudy (anonymous):

ok so im clueless right now sorry i couldnt help

OpenStudy (anonymous):

its fine...XD

OpenStudy (asad786):

alright let me work through this

OpenStudy (karolynxx):

It is B , (2,-1)

OpenStudy (karolynxx):

what else?

OpenStudy (anonymous):

k thanks, I will post the next one :D

OpenStudy (asad786):

yup b

OpenStudy (karolynxx):

sorry i meant (-2,1)

OpenStudy (asad786):

ask me could you medal me

OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? x-4y=6 3x+y=-8 A. (2, −1) B. (1, −9) C. (−2, −2) D. (−6, −3)

OpenStudy (anonymous):

@karolynxx ...you left me... XD oh well...

OpenStudy (anonymous):

...can anyone help me

OpenStudy (karolynxx):

im still here lol

OpenStudy (anonymous):

XD

OpenStudy (karolynxx):

i believe C

OpenStudy (anonymous):

ok :D

OpenStudy (anonymous):

can I post the next one?

OpenStudy (karolynxx):

yes

OpenStudy (anonymous):

k, just a sec XD

OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? -3x+4y=-20 x+4y=-4 A. (8, 1) B. (−12, 2) C. (−4, −8) D. (4, −2)

OpenStudy (anonymous):

I really dont know this stuff.... :(

OpenStudy (anonymous):

im ganna close this question and re-post it XD